40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3257     Entry time: Wed Jul 21 12:20:29 2010
Author: josephb, kiwamu 
Type: Update 
Category: Computers 
Subject: Megatron temporarily disconnected, c1iscex firewalled, green FE test 

We are moving towards a first test of getting Kiwamu's green locking signals into the new front end at the new X end, as well as sending signal out to the green laser temperature control.

Towards that end, we borrowed the router which we were using as a firewall for megatron.   At the moment, megatron is not connected to the network.  The router (a linksys N wire router), was moved to the new X end, and setup to act as a firewall for the c1iscex machine.

At this point, we need to figure which channels of the DAC correspond to which outputs of the anti-imaging board (D000186) and coil driver outputs.  Ideally, we'd like to simply take a spare output from that board and bring it to the laser temperature control.  The watchdogs will be disabled when testing to avoid any unfortunate mis-sent signals to the coils.  It looks like it should be something like channels 6,7,8 are free, although I'm not positive if thats the correct mapping or if there's a n*8 + 6,7,8 mapping.

The ADC should be much easier to determine,  since we only have a single 16 channel set coming from the lemo breakout box.  Once we've determined channels, we should be all set to do a test with the green system.

ELOG V3.1.3-